Your father offers you a choice of $110,000 in 12 years or $47,500 today. Use Appendix B as an approximate answer, but calculate your final answer using the formula and financial calculator methods.

a-1. If money is discounted at 10 percent, what is the present value of the $110,000? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
  



a-2. Which offer should you choose?
  

  • $110,000 in 12 years

  • $47,500 today


b-1. Now assume the offer is $110,000 in 9 years or $47,500 today. What is the present value of the $110,000 at 10 percent for 9 years? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
  



b-2. Now, which offer should you choose?
  

  • $47,500 today

  • $110,000 in 9 years
Answer #1

a-1.

present value of the $110,000 = 110000/(1+10%)^12 = 35049.38995 = 35049.39 (Rounded to 2 decimal places)

a-2.

$47500 today should be chosen as it is more than present value of $110,000 in 12 years which is 35049.39

b-1.

present value of the $110,000 for 9 years => 110000/(1+10%)^9 = 46650.73802 = 46650.74 (Rounded to 2 decimal places)

b-2.

$47500 today should be chosen as it is more than present value of $110,000 in 9 years which is 46650.74